Suppose you have a house worth $200,000 (wealth). Your utility of wealth is given by U(w) = ln(w). There is a small chance that a fire will damage your house causing a loss of $75,000. You estimate there is a 2% chance of fire.

a) What is your expected wealth?

b) What is your expected utility from owning the house?

c) Suppose you can add a fire detection/prevention system to your house. This would reduce the chance of a bad event to 0 but it would cost you $C to install. What is the most you are willing to pay for the security system? (Here is an identity you will find useful 


1
Expert's answer
2021-08-19T12:23:59-0400

Given,

Utility function:

"U(w)=Ln(w)"

(a). The formula for expected wealth is given below:

Expected wealth=probability of fire×wealth in case of fire+probability of

 no fire×wealth in case of no fire

"Expected \space wealth=2100×(200,000−75,000)+\frac{98}{100}×200,000\\Expected\space wealth=2,500+196,000\\Expected \space wealth=\$198,500"

The expected wealth is $198,500.


(b). The value of expected utility is given below:

"Expected\space utility=\frac{2}{100}ln(125,000)+\frac{98}{100}ln(200,000)\\Expected\space utility=\frac{2}{100}×11.7361+\frac{98}{100}×12.2061\\Expected \space utility=0.2347+11.9620\\Expected\space utility=12.1967"

The value of expected utility is 12.1967 approx.


(c). The maximum amount that can be paid is the difference between the certainty equivalent (CE) and the current value of wealth.

"U(w)=expected\space utility\\ln(w)=12.1967\\w=e^{12.1967}w=198,134.2282\\Maximum \space payment=200,000−198,134.2282\\Maximum \space payment=\$1865.77 approx"

The most that can be paid is $1865.77 approx.